# Tackling Cyberattacks through AI-based Reactive Systems: A Holistic Review and Future Vision

### v1

@@ V Comparative Exploration of Analyzed Works
[81] 2018
Fault tree Analysis
Fuzzy Theory
Vulnerability Assesment
Multiple
[85] 2018
Attack Graphs
Mathematical Model
Minimising Risk
Private Network
[89] 2019
Attack Graphs redefined
A* algorithm
Minimising Risk
Computing time
Multiple
[92] 2019
Machine Learning
Neural Network - Self-taught Learning
Threat Detection
Variety of Classification
Multiple
[96] 2019
Bayesian Network
Pareto Optimal Solutions
Minimising Risk
Countermeasure Selection Speed
Critical Infrastructure
[97] 2020
Hypergraphs
StRORI
Minimising Risk and Cost
Countermeasure Selection Speed/Estimated Risk/Losses
Multiple
[99] 2020
Attack Graphs
Attack Defense Trees
Minimising Risk
Countermeasure Selection Speed
Multiple
[101] 2020
Bayesian Network
Probabilistic Model
Minimising Risk
Implementation Cost and Impact
Multiple
[102] 2020
Dynamic Countermeasure Knowledge
Countermeasure Standard Proposal
ROI Score
Multiple
[104] 2020
Machine Learning
Evolution Algorithm - Reinforcement Learning (Markov Decision Process)
Minimising Risk
Positive Rate
Microservices Architecture
[107] 2020
Hierarchical Risk Correlation Tree
Autonomous Response Controller
Minimising Risk
Countermeasure Selection Speed
Cyber–Physical Systems
[108] 2021
Machine Learning
Artificial Immune Systems
Minimising Risk
Countermeasure Selection Speed
Multiple
[110] 2021
Machine Learning
Evolution Algorithm - Reinforcement Learning
Minimising Risk
Positive Rate
Elastic Applications
[112] 2022
Game Theory
Nash Equilibrium
Minimising Risk and Cost
Private Network
[117] 2022
Attack Graphs
Mathematical Model
Minimising Risk
Countermeasure Selection Speed
Critical Infrastructure
[119] 2022
Attack Graphs
Evolution Algorithm - Reinforcement Learning
Minimising Risk and Cost
Positive Rate
Industrial Control Systems
[120] 2022
Machine Learning
Deep Learning - Supervised Learning
Threat Detection
Positive Rate
Internet of Things
[123] 2022
Attack Graphs
Evolution Algorithm - Reinforcement Learning
Threat Detection
Positive Rate
Multiple
[125] 2022
Machine Learning
Deep Learning
Threat Detection
Positive rate
Critical Infrastructure
[126] 2022
Machine learning
Evolution Algorithm - Reinforcement Learning
Mitigate Attack
Positive Rate
Multiple
[128] 2023
Attack Trees
Heuristic - Probabilistic methods
Calculation of Cost-damage probabilities in Attacks
Calculation Speed
Multiple
TABLE I: Features of analyzed papers
@@ V-B Algorithm
This diversification in the amount of algorithms may negatively affect the development of an autonomous reaction system. That is because it may prevent researchers from following a common strategy. Furthermore, this variety could present various problems when integrating components into organisations or make it harder to compare the various proposals due to their differing similarities.
@@ V-C Strategy Objectives
In the case of the “Strategy Objectives” column, it is readily observable that the most frequent objective is risk minimization. This is achieved through the selection or generation of countermeasures to counteract threats. The risk minimization objective can also include values in which the costs associated with implementing countermeasures are taken into account. These reflect the goals and needs perceived by researchers in the field, where the objective is to optimize cybersecurity investment without compromising the protection of critical systems and assets. Additionally, there is an effort to strike a balance between mitigating identified threats and the resources available at that moment.
@@ V-D Measurements and results
Additionally, it is worth noting the use of metrics such as computation time or countermeasure selection time. These metrics shed light on the efficiency and speed at which the countermeasures are identified and implemented. This efficiency is particularly relevant in environments that face threats, where rapid response times can make a significant difference in containing potential damages and minimizing risks.
To conclude the analysis of this column, it may be interesting to note that during the survey, many risk measures could be identified in terms of vulnerabilities, such as CVSS or CVE, while for countermeasures no value has been found that could classify them.
@@ V-E Use Case
When examining the “Use Case” column, it becomes evident that the value “All” holds significant prominence. This indicates that most of the works analyzed in this survey do not tailor their solutions to a specific system. Consequently, approaches marked with this value possess the advantage of adaptable deployment across diverse environments, ensuring highly efficient protection for most scenarios. However, it is important to consider that by not tailoring the proposal to a specific system, certain specific areas within different environments may remain unprotected. For instance, in critical infrastructure environments, meticulous consideration of all details is imperative, given the heightened significance of these devices compared to common devices.
@@ VI-A Unified Frameworks
One of the primary research challenges lies in establishing standardized procedures and unified structures within AI-powered threat response systems. The current challenge with these systems is caused by the lack of a common baseline. This lack of shared foundations can have several significant implications and consequences. For instance, absence of interoperability between tools or systems that have to be coordinated for the exchange of information. Alternatively, it may make it difficult to compare different proposals because they have no common basis for benchmarking.
By developing unified frameworks that provide a shared basis for the integration of various AI techniques, consistency and interoperability could be introduced into these response systems. Such an approach could accelerate development processes and pave the way for collaborative advances that could strengthen cohesion among researchers and practitioners.
@@ VI-B Collaborative AI
To improve collaboration between different AI methodologies, researchers could focus on creating a standardized platform that facilitates the test of the integration of different AI techniques. In this case, different methodologies could be assessed and ranked according to their advantages and weaknesses based on specific tasks. Subsequently, complex systems could be designed that unify models according to their strength, thus, building hybrid models that design the reaction problems and assign tasks to different AIs.
@@ VI-C Evolutive AI
Adaptation and evolution mechanisms of AI models used in threat response systems represent another interesting research perspective. Due to the dynamic nature of cyber threats, AI models require evolution and adaptation. This is because threats are evolving, and new and unknown vulnerabilities are being discovered regularly. Then, similarly, as threats evolve, AI models must be dynamically updated. This challenges presents an intriguing research trajectory to facilitate real-time updates driven by the evolving threat landscape. The real-time evolution would allow AI models to remain effective over time and could avoid the obsolescence that some of the models in this survey may exhibit.
It is important to note that while AI can be designed to be evolutive, it can also present a series of potential issues. Among these is the need to re-train models, which could present problems of time and loss of prior knowledge. Moreover, this re-training would have to be programmed, and this presents another dilemma in terms of determining the optimal time for this evolution.
To address this challenge, automated mechanisms should be designed to update AI models. The implementation of a feedback loop between real-world threat data and AI models can ensure regular and timely updates. Besides, cooperation with threat intelligence providers can be established to provide such threat data. This would allow access to updated information for adapting the models. Alternatively, techniques such as transfer learning and RL [119, 135, 104] can be applied to facilitate faster adaptation of AI models to new threat scenarios.
@@ VI-D Collaborative Knowledge Exchange
To achieve effective cybersecurity, collaboration is a strong foundation. Collaboration allows for an approach that promotes research towards collaborative learning. By enabling organizations and states to collaborate in sharing information/knowledge for training AI models without compromising data privacy, a collective defense mechanism could be developed and, due to the joint effort, become potentially more effective. Particularly, this line of research could trigger the generation of shared knowledge leading to a more robust and informed defence against evolving threats.
To address this challenge, the first step should be to establish trusted frameworks and protocols for secure collaboration. To share threat data securely without revealing sensitive information. Next, efforts should be made to find an effective way of sharing information that reaches widely and equally. For example, a centralized platform that could act as a threat intelligence intermediary.
@@ VI-E AI and Human Experts
In the context of collaboration is not limited to the organizational environment. Indeed, effective collaboration between humans and AI systems presents another intriguing scenario. Future research could focus on designing intuitive user interfaces that facilitate interaction between human experts and AI-based threat response systems. This collaboration empowers decision-makers by providing them with AI-generated knowledge while allowing for human oversight and judgment, which improves the overall threat response process. In addition, this also addresses the need to explain the decisions made by autonomous response systems, named as eXplainable Artificial Intelligence (XAI) [138], requiring the technicians operating these systems to have control over the decisions being made.
@@ VI-F Countermeasure standardization
To effectively standardize countermeasures, efforts should be dedicated to the categorization of threats and vulnerabilities in a standardized way. Developing a common language for countermeasures that fits this categorization can help standardize responses [108, 102]. In addition, leveraging ML and natural language processing to analyze and classify emerging threats can automate the generation of countermeasures.
To effectively standardize countermeasures, little efforts should be dedicated to the categorization of threats and vulnerabilities in a standardized way. Reduced efforts can be attributed to the fact that, as mentioned above, there are standards for the representation of vulnerabilities and threats that are widely used, such as CVE and CVSS. This implies that progress should continue to be made in this field, however
it could be interesting for academia to research a method for standardization of countermeasures that provides a scoring. In the same way as the categorization of vulnerabilities, such a score could be used to assess the effectiveness of a countermeasure. Thus helping the optimal selection of countermeasures by a reaction system. In essence, developing a common language for countermeasures that fits this categorization can also help standardize responses [108, 102].
@@ VI-H Summary
To enhance the reader’s understanding, an illustrative support, represented by Figure 7, has been incorporated. This figure visually presents the research challenges, illustrating their interconnections through arrows. In particular, an additional parameter called “Collaboration” has been introduced to indicate whether a research pathway requires stakeholder collaboration.
Figure 7: Graph with research challenge interconnections

### v2

@@ V Comparative Exploration of Analyzed Works
[83] 2018 Fault tree Analysis Fuzzy Theory Vulnerability Assessment - Multiple
[87] 2018 Attack Graphs Mathematical Model Minimizing Risk - Private Network
[91] 2019 Attack Graphs redefined A* algorithm Minimizing Risk Computing time Multiple
[94] 2019 Machine Learning Neural Network - Self-taught Learning Threat Detection Variety of Classification Multiple
[98] 2019 Bayesian Network Pareto Optimal Solutions Minimizing Risk Countermeasure Selection Speed Critical Infrastructure
[99] 2020 Hypergraphs StRORI Minimizing Risk and Cost Countermeasure Selection Speed/Estimated Risk/Losses Multiple
[101] 2020 Attack Graphs Attack Defense Trees Minimizing Risk Countermeasure Selection Speed Multiple
[103] 2020 Bayesian Network Probabilistic Model Minimizing Risk Implementation Cost and Impact Multiple
[104] 2020 Dynamic Countermeasure Knowledge - Countermeasure Standard Proposal ROI Score Multiple
[106] 2020 Machine Learning Evolution Algorithm - Reinforcement Learning (Markov Decision Process) Minimizing Risk Positive Rate Microservices Architecture
[109] 2020 Hierarchical Risk Correlation Tree Autonomous Response Controller Minimizing Risk Countermeasure Selection Speed Cyber–Physical Systems
[110] 2021 Machine Learning Artificial Immune Systems Minimizing Risk Countermeasure Selection Speed Multiple
[112] 2021 Machine Learning Evolution Algorithm - Reinforcement Learning Minimizing Risk Positive Rate Elastic Applications
[114] 2022 Game Theory Nash Equilibrium Minimizing Risk and Cost - Private Network
[119] 2022 Attack Graphs Mathematical Model Minimizing Risk Countermeasure Selection Speed Critical Infrastructure
[121] 2022 Attack Graphs Evolution Algorithm - Reinforcement Learning Minimizing Risk and Cost Positive Rate Industrial Control Systems
[122] 2022 Machine Learning Deep Learning - Supervised Learning Threat Detection Positive Rate Internet of Things
[125] 2022 Attack Graphs Evolution Algorithm - Reinforcement Learning Threat Detection Positive Rate Multiple
[127] 2022 Machine Learning Deep Learning Threat Detection Positive rate Critical Infrastructure
[128] 2022 Machine learning Evolution Algorithm - Reinforcement Learning Mitigate Attack Positive Rate Multiple
[130] 2023 Attack Trees Heuristic - Probabilistic methods Calculation of Cost-damage probabilities in Attacks Calculation Speed Multiple
TABLE I: Side-by-side comparison of the selected works based on the features described in Section III
@@ V-A Methodology
From a practical point of view, one can say that it makes sense to employ machine learning-based methodologies, especially in enterprise environments. In these enterprise networks, where numerous vulnerabilities may exist, a large amount of data is constantly being generated. Each of these points can provide a vast amount of information that needs to be processed and analyzed to provide effective countermeasures against threats detected. ML therefore can extract meaningful patterns and make accurate predictions from these large datasets. It could proactively analyze threat alerts generated by IDSs and seek appropriate countermeasures. Furthermore, ML has the advantage of being able to adapt and continuously improve over time as more information is provided.
In the application of graph-based methodologies, within organizational settings such as corporate or military environments, several significant benefits for risk management and security can be identified. These methodologies can provide a clear visual representation of potential attack routes that could compromise security. This allows security teams to quickly understand potential threats and infrastructure hotspots. Similarly, it allows for the representation of multiple attack scenarios, providing a more complete understanding of potential threats and their possible impacts on the organization.
@@ V-B Algorithm
This diversification in the amount of algorithms may negatively affect the development of an autonomous reaction system. That is because it may prevent researchers from following a common strategy. Furthermore, this variety could present various problems when integrating components into organizations or make it harder to compare the various proposals due to their differing similarities.
From a pragmatic standpoint, the diversity of algorithms found in the papers analyzed can be attributed to several reasons. On the one hand, different problems and research contexts require specific algorithmic approaches to solve them. For example, while data classification problems may benefit from DL algorithms such as NN, optimization problems may require evolutionary or game theory-based algorithms. The variety of algorithms may reflect the diversity of tools available in the fields of data science and AI. Researchers can select the most appropriate algorithm according to their prior knowledge, experience, and the nature of the problem they are tackling. This diversity of tools may result in the choice of different algorithmic approaches to similar problems, depending on individual preferences and computational resource constraints.
This diversity reflects the multifaceted and constantly changing nature of these fields, where innovation and experimentation are fundamental to the advancement of knowledge and technology, and where the lack of research in the field of reactive systems has not laid a foundation for the creation of robust reactive systems that work in an enterprise environment.
@@ V-C Strategy Objectives
In the case of the “Strategy Objectives” column, it is readily observable that the most frequent objective is risk minimization. This is achieved through the selection or generation of countermeasures to counteract threats. However, it is important to note that risk minimization can include values in which the costs associated with implementing countermeasures are taken into account. These reflects the goals and needs perceived by researchers in the field, where the objective is to optimize cybersecurity investment without compromising the protection of critical systems and assets. Nevertheless , it is worth questioning whether this pursuit of cost minimization could lead to compromises in the effectiveness of countermeasures and ultimately in system security.
Furthermore , the usefulness of AI in this case lies in its adaptability to the strategy chosen for each case. The application of AI methods, such as machine learning models, may vary depending on the specific threat and the organization’s security objectives. Ultimately, integrating A into cyber security strategies could not only improve response capabilities, but also facilitate a more flexible response to evolving cyber threats depending on the strategy the organization chooses to pursue.
Additionally, there is an effort to strike a balance between mitigating identified threats and the resources available at that moment. This search for efficiency in resource allocation is essential to ensure an effective response to cyber threats, but it may also pose significant challenges. For example, it could generate tensions between the need to implement rapid countermeasures and the limited availability of financial or technical resources.
@@ V-D Measurements and results
Moreover, the emphasis on metrics such as computation time and countermeasure selection time underscores the importance placed on efficiency and speed in the context of response systems. While rapid response times are indeed crucial for mitigating potential threats, an overemphasis on speed may inadvertently overshadow other critical factors, such as the accuracy and efficacy of the selected countermeasures. This drawback raises concerns about whether the evaluation criteria adequately capture the multifaceted nature of cybersecurity challenges and response strategies.
In this context, AI integration could offer competitive advantages. AI systems could help in streamlining automation tasks by taking into account the different ways of measuring the results that a system could have. Trying to adjust different system configurations to get better results depending on the metrics used.
The absence of comprehensive risk measures for countermeasures is a notable gap in the analysis. While vulnerabilities can be assessed using established measures such as CVSS or CWSS, a corresponding framework for assessing countermeasures seems to be missing. This shortcoming undermines the ability to effectively compare and prioritize countermeasures based on their potential impact and their effectiveness in mitigating specific threats. It also highlights a more general problem in the domain, where the focus on identifying vulnerabilities or exposures often exceeds the focus on developing and evaluating robust countermeasures.
In conclusion, while the analysis provides valuable insights into certain aspects of performance and efficiency, it also reveals underlying limitations and areas for improvement. A more balanced and comprehensive approach to performance assessment, covering a wider range of parameters and taking into account the effectiveness of countermeasures to mitigate various threats, is essential to move forward.
@@ V-E Use Case
When examining the “Use Case” column, it becomes evident that the value “All” holds significant prominence. This indicates that most of the works analyzed in this survey do not tailor their solutions to a specific system. While this approach offers the advantage of versatile deployment in a variety of contexts, allowing for effective protection in a range of scenarios, it also raises critical questions about the suitability of these one-size-fits-all solutions. By opting for a generic approach, there is a risk of overlooking the complex and nuanced requirements and vulnerabilities inherent in different systems and environments. This lack of customization may leave certain areas within various environments inadequately protected, especially in high-risk sectors such as critical infrastructure. In these environments, where the consequences of a cyber attack can be catastrophic, a more tailored and meticulous approach to cyber security may be required.
Furthermore, the prevalence of the term “All” may indicate a tendency to oversimplify or generalize in the development and evaluation of cyber security solutions. While adaptive deployment is undoubtedly valuable, it must be accompanied by a thorough understanding of the specific risks and challenges posed by different use case scenarios. This could lead to a false sense of security and leave critical systems vulnerable.
In conclusion, while the preference for the designation “All” in the “Use case” column highlights the adaptability of the methodologies or systems analyzed, it also underlines the need for a more critical and tailored approach to the development of cybersecurity solutions.
@@ VI-A Unified Frameworks
One of the primary research challenges lies in establishing standardized procedures and unified structures within AI-powered threat response systems. The current challenge with these systems is caused by the lack of a common baseline. This lack of shared foundations can have several significant implications and consequences.
The absence of a shared baseline results in a significant lack of interoperability between tools and systems, preventing a potential fluent exchange of information between researchers and security teams in practical scenarios. . This fragmentation poses significant challenges when comparing different proposals, as there is no standardized benchmark for assessment.
To overcome this obstacle, it is imperative to develop unified frameworks. Such frameworks would outline standardized protocols and structures for the seamless integration of various AI techniques, thereby promoting consistency and interoperability. The implementation of such frameworks would streamline the development process and foster collaboration between researchers and practitioners.
@@ VI-B Collaborative AI
To improve collaboration between different AI methodologies, researchers could focus on creating a standardized platform that facilitates the test of the integration of different AI techniques. Through this platform, different methodologies could be evaluated and ranked according to their effectiveness for specific tasks, taking into account their respective strengths and weaknesses. Subsequently, sophisticated systems could be designed that unify models according to their strength, thus, creating hybrid models that design the reaction problems and assign tasks to different AIs.
@@ VI-C Evolutive AI
Adaptation and evolution mechanisms of AI models used in threat response systems represent another interesting research perspective. Due to the dynamic nature of cyber threats, AI models require evolution and adaptation. This is because threats are evolving, and new and unknown vulnerabilities are being discovered regularly. This requires real-time adaptation of AI models to remain effective over time and to avoid possible obsolescence.This challenges presents an intriguing research trajectory to facilitate real-time updates driven by the evolving threat landscape. The real-time evolution would allow AI models to remain effective over time and could avoid the obsolescence that some of the models in this survey may exhibit.
Implementing evolutive AI presents practical challenges, particularly concerning re-training models. Re-training can be time-intensive and may result in the loss of prior knowledge, impacting operational efficiency. Additionally, determining the optimal timing for model updates is essential to ensure that systems are responsive to emerging threats without disrupting ongoing operations. To address these challenges practically, automated mechanisms for model updates are crucial. Establishing a robust feedback loop between real-world threat data and AI models enables regular and timely updates.
Furthermore, practical techniques such as transfer learning and RL [121, 137, 106] can expedite the adaptation process. Transfer learning allows models to leverage knowledge from pre-trained models, reducing the need for extensive re-training from scratch. RL, on the other hand, can simulate various threat scenarios and enable AI models to learn and adapt in real-time based on feedback from the environment.
Collaboration with threat intelligence providers is another practical strategy. Partnering with these providers can provide direct access to the latest threat information, enabling faster and more informed decision-making for model adaptation. By integrating real-time threat feeds into AI systems, researchers can ensure that models are updated with the most current threat data.
@@ VI-D Collaborative Knowledge Exchange
To achieve effective cybersecurity, collaboration is a strong foundation. Collaboration allows for an approach that promotes research towards collaborative learning. By enabling organizations and states to collaborate in sharing information/knowledge for training AI models without compromising data privacy, a collective defense mechanism could be developed and, due to the joint effort, become potentially more effective. Particularly, this line of research could trigger the generation of shared knowledge leading to a more robust and informed defense against evolving threats.
To address this challenge, the initial step should be to establish trusted frameworks and protocols for secure collaboration, ensuring the sharing of threat data without compromising sensitive information. Subsequently , efforts should be made to find an effective way of sharing information that reaches widely and equally. For example, a centralized platform that could act as a threat intelligence intermediary.
@@ VI-E AI and Human Experts
In the context of collaboration is not limited to the organizational environment. Indeed, effective collaboration between humans and AI systems presents another intriguing scenario. Future research could focus on designing intuitive user interfaces that facilitate interaction between human experts and AI-based threat response systems. This collaboration empowers decision-makers by providing them with AI-generated knowledge while allowing for human oversight and judgment, which improves the overall threat response process.
A practical approach to confront this challenge involves designing user interfaces that prioritize usability and user experience. Researchers can conduct usability studies and apply human-centered design principles to create intuitive interfaces that foster cooperation between humans and AI. An essential aspect of this collaboration is the integration of eXplainable Artificial Intelligence (XAI) techniques [140]. XAI enables AI systems to explain their decisions in a transparent and understandable manner, allowing human experts to validate and interpret AI-generated insights.
@@ VI-F Countermeasure standardization
To effectively standardize countermeasures, efforts should be dedicated to the categorization of threats and vulnerabilities in a standardized way. Developing a common language for countermeasures that fits this categorization can help standardize responses [110, 104]. By establishing a common language for countermeasures that fits this categorization, responses can be standardized across organizations and industries. This not only facilitates communication and collaboration, but also ensures consistency and interoperability in the application of security measures.
In addition, the integration of advanced technologies such as ML and natural language processing can significantly improve the effectiveness of countermeasure standardization efforts. ML algorithms can analyses large amounts of threat data to identify patterns and trends, enabling automatic classification of emerging threats. Similarly, NLP techniques can extract valuable information from unstructured data sources such as security reports and threat intelligence sources, further enriching the threat analysis process. By leveraging these technologies, organizations can streamline the generation and deployment of countermeasures.
However, while there are existing standards such as CVE and CVSS for vulnerability representation, progress should continue in this field. It could be beneficial for academia to research a method for standardization of countermeasures that provides a scoring. In the same way as the categorization of vulnerabilities, such a score could be used to assess the effectiveness of a countermeasure. Thus helping the optimal selection of countermeasures by a reaction system. In essence, developing a common language for countermeasures that fits this categorization can also help standardize responses [110, 104].
@@ VI-G Generative AI
A significant aspect of this challenge lies in the integration of generative AI models into existing cyber security frameworks. Unlike traditional reactive systems, which rely on predefined rules and signatures to detect and mitigate threats, generative AI models have the ability to autonomously generate new attack scenarios and vulnerabilities. This dynamic nature introduces complexities in threat detection and response.
Furthermore, the ethical implications of using generative AI models in cybersecurity cannot be overlooked. The use of AI-generated adversarial examples, for example, raises concerns about the potential for unintended damage or exploitation. Investigating the ethical considerations and implications of integrating generative AI models into reactive systems is therefore crucial to ensure responsible and equitable cybersecurity practices.
@@ VI-H Summary
To enhance the reader’s understanding, an illustrative support, represented by Figure 9, has been incorporated. This figure visually presents the research challenges, illustrating their interconnections through arrows. In particular, an additional parameter called “Collaboration” has been introduced to indicate whether a research pathway requires stakeholder collaboration. The introduction of “Collaboration” underscores the importance of cooperative efforts in addressing these complex and multifaceted challenges.
For instance, developing “Collaborative AI” requires AI systems to work together and share information. This inherently requires collaboration among different AI developers, researchers, and possibly end-users to ensure that these systems can interact and learn from each other efficiently. Similarly, “Collaborative Knowledge Exchange” involves the continuous flow of information and insights between AI systems and human experts, requiring a collaborative framework to facilitate this exchange.
A central element in this network of research challenges is the “Unified Frameworks” challenge. This challenge is pivotal as it connects with “Collaborative Knowledge Exchange” and “Collaborative AI,” emphasizing the need for comprehensive frameworks that can integrate different AI systems and facilitate seamless interactions.
“Countermeasure Standardization” challenge also demands collaboration as it involves creating standardized approaches to mitigate risks associated with AI systems.“AI and Human Experts” focuses on the integration and synergy between AI systems and human expertise. Lastly, “Evolutive AI” stands as an independent challenge that does not directly require collaboration with other research pathways. However, its development can still benefit from collaborative insights, particularly in understanding how AI systems can evolve and adapt over time.
Figure 9: Graph showing the interconnections of the research challenges according to their similarity and their dependence on collaboration.